OpenStudy (anonymous):

Two cars are at rest, facing each other and 200m apart. If Car 1 accelerates at 6.00m/s^2 up to maximum speed of 30.0m/s and Car 2 accelerates at 5.00m/s^2 up to a maximum speed of 35.0m/s, how long does it take he cars to meet?

OpenStudy (anonymous):

its a problem of kinematics

OpenStudy (anonymous):

Time to for car 1 to reach max velocity is 5 second and 7 seconds for car 2

OpenStudy (anonymous):

then i used d=vo+1/2at^2 to find the distance for the cars at 5 and 7 seconds and i got for 5 seconds car 1 and 2 are 75 and 62.5. For 7 seconds i got 147 for car 1 and 122.5 for car 2

OpenStudy (anonymous):

all i know is the collision point is between 5 and 7 seconds, im stuck there

OpenStudy (anonymous):

for first car caluclate distance taking initial speed as 0m/s and final speed as 30m/s and acceleration as 6m/s^2

OpenStudy (anonymous):

then in similar way caluclate for second

OpenStudy (anonymous):

for car 1 i got D 75m, T 5s, and car 2 i got D 122.5m, T 7 s

OpenStudy (anonymous):

hey@bukloon ur half way now take the cars acceleration as 0m/sec^2

OpenStudy (anonymous):

you mean make acceleration for both cars and make them 0m/sec^2?

OpenStudy (anonymous):

total distance will be 200-(75+122.5) and then s=(u1*t+2)+u2*t) 2.5=(30*t+2)+(35*t) then caluclate it for t t+7 wud be the answer

OpenStudy (anonymous):

ha ya @BukLoon b'cause in question its mentioned that the max. speed so after acquiring that speed there wud be no acceleration

OpenStudy (anonymous):

they wud b moving vth a constant speed

OpenStudy (anonymous):

I cant picture they both stop acceleration once they reach their constant speed

OpenStudy (anonymous):

can**

OpenStudy (anonymous):

answer is 6.08 s

OpenStudy (anonymous):

hmm so i got somewhere wrong stop i will check it out

OpenStudy (anonymous):

k the first car accelerated for 5 sec and in nxt two sec its travellling its travelling 60mts so problem is here

OpenStudy (anonymous):

so now caluclate distaance for 5sec for second car

OpenStudy (anonymous):

75+60

OpenStudy (anonymous):

noo i will make u understand

OpenStudy (anonymous):

in first 5 sec first car reached max speed that is 30m/sec bt second in first 5 sec didn't reach its max speed and its accelerating and it travelled 62.5mts

OpenStudy (anonymous):

did u get that till now

OpenStudy (anonymous):

yah im at that part, so at 7 seconds car one would be going 135 m and car 2 would be going 122.5

OpenStudy (anonymous):

hmm that is beyond 200mts

OpenStudy (anonymous):

so now we must apply uniform motion for first car after 5 sec and accelerated mation for second car

OpenStudy (anonymous):

so now s=200-(75+62.5) s=u1*t+(u2*t+1/2*a*t^2) so, 62.5=30*t+(25*t+1/2*5*t^2)

OpenStudy (anonymous):

did u get it.......@BukLoon

OpenStudy (anonymous):

yes its worked out, thanks you Ravi!

OpenStudy (anonymous):

mention not .............@ BukLoon

OpenStudy (anonymous):

did u understand...........

OpenStudy (anonymous):

d=vt+(vit+1/2at^2)

ganeshie8 (ganeshie8):